Santana Mortgage Company uses a process cost system to accumulate costs in its Application Department. When an application is completed, it is forwarded to the Loan Department for final processing. The following processing and cost data pertain to September.

1. Applications in process on September 1, 100
2. Applications started in September, 900
3. Completed applications during September, 700
4. Applications still in process at September 30 were 100% complete as to materials (forms) and 60% complete as to conversion costs.
Beginning WIP:
    Direct materials $1,320
    Conversion costs 4,230
September costs:
    Direct materials $3,880
    Direct labor 12,210
    Overhead 6,968


Materials are the forms used in the application process, and these costs are incurred at the beginning of the process. Conversion costs are incurred uniformly during the process.

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Determine the equivalent units of service (production) for materials and conversion costs.

Materials

Conversion Costs

The equivalent units of service (production)

  

  

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Compute the unit cost. (Round unit costs to 2 decimal places, e.g. 2.25.)

Materials

Conversion Costs

Unit costs

$

$

  

  

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Prepare a cost reconciliation schedule. (Round unit costs to 2 decimal places, e.g. 2.25 and final answers to 0 decimal places, e.g. 1,225.)

Costs accounted for:

Transferred out

$

Work in process, September 30

   Materials

$

   Conversion costs

Total costs

$
Answer #1

Solution

Material Conversion
Equivalent Units of Production                                  1,000                         880

.

Material Conversion
Cost per equivalent Unit $ 5.20 $ 26.60

.

Costs accounted for:
    Transferred out $            22,260
    Work in process, September 30
       Materials $           1,560
       Conversion costs $           4,788 $              6,348
    Total costs $            28,608

Working

Reconciliation of Units
A Beginning WIP                         100
B Introduced                         900
C=A+B TOTAL                     1,000
D Transferred out                         700
E=C-D Ending WIP                         300

.

Statement of Equivalent Units(Weighted average)
Material Conversion
Units Complete % Equivalent units Complete % Equivalent units
Transferred out                                      700 100%                          700 100%                         700
Ending WIP                                      300 100%                          300 60%                         180
Total                                  1,000 Total                       1,000 Total                         880

.

Cost per Equivalent Units (Weighted average)
COST Material Conversion TOTAL
Beginning WIP Inventory Cost (A) $ 1,320 $ 4,230 $ 5,550
Cost incurred during period (B) $ 3,880 $ 19,178 $ 23,058
Total Cost to be accounted for (C=A+B) $ 5,200 $ 23,408 $ 28,608
Total Equivalent Units(D)                     1,000                           880
Cost per Equivalent Units (E=C/D) $ 5.20 $ 26.60 $ 31.80

..

Statement of cost (Weighted average)
Cost Equivalent Cost/unit Ending WIP Transferred
Units Cost Allocated Units Cost Allocated
Material $ 5.20                         300 $             1,560.00                             700 $            3,640.00
Conversion $ 26.60                         180 $             4,788.00                             700 $          18,620.00
TOTAL $ 28,608 TOTAL $ 6,348 TOTAL $ 22,260